Draw a number line to graph the number:3.2;-0.7;-5;3/2;3/4;-5/2
1
Expert's answer
2020-10-08T16:53:49-0400



A(3.2), B(-0.7) C(-5), D("\frac{3}{2}") E("\frac{3}{4}") F(-"\frac{5}{2}")